OpenStudy (anonymous):

Minimize S = x + 19y with xy = 19 and both x and y > 0. and also Minimize 8x^2+8y^2 with xy^2=16 please help! :(

OpenStudy (anonymous):

alrigt the first problem... you have sum=x+19y and xy=19

OpenStudy (anonymous):

use the second equation to solve for x or y since x is by itself i'd solve for x y=19/x

OpenStudy (anonymous):

x=19/y i mean next put this into the first equation S=19/y+19y

OpenStudy (anonymous):

next take S' \[S'=-19y^{-2}+19\]

OpenStudy (anonymous):

thank you!